# Unsupervised Multimodal Graph-based Model for Geo-social Analysis

### v1

@@ 5.1 Topic Interpretation
Other topics represent broader or more indirect engagement with the event. Topic 8 in MonoGraph and Topic 5 and in MultiGraph, captured meteorological observations, with keywords such as “temperature,” “humidity,” “clouds,” “rain,” and “wind,” are more evenly distributed across central Texas, including Austin and San Antonio. These spatial patterns suggest contributions from users outside the most severely affected areas, often in the form of commentary or weather updates rather than first-hand impact reports. Weather-related discourse is further reflected in Topics 2, 7, 5 and 9, in MonoGraph and Topics 3, 4, 9, 6 and 8 in MultiGraph, featuring overlapping terms such as “pressure,” “gust,” “humidity,” “wind,” and “clouds.” While this lexical overlap reduces topic distinctiveness, it underscores the recurring presence of weather-related signals across the corpus, a common characteristic in disaster-related social media data. This also illustrates the trade-off of fixing the number of topics ($K=10$), which facilitates stability but can lead to partial redundancy. Finally, content-sharing and localized observational behavior emerge as a consistent pattern. For example, certain clusters contain terms such as “posted,” “photo,” and place names like “meyerland,” “katy,” and “downtown,” suggesting the dissemination of real-time visual documentation and location-tagged updates by affected residents. This pattern reflects well-known crisis communication behavior on social media, where users share on-the-ground images to inform others and raise awareness.
@@ 6 Discussion and Limitations
Across all datasets, both MonoGraph and MultiGraph consistently outperformed baseline methods, achieving higher intra-cluster cohesion and lower inter-cluster similarity. Their relative strengths varied with event characteristics, suggesting sensitivity to spatial scale, discourse patterns, and data quality. Fixing the number of clusters at $K=10$ facilitated direct cross-model comparisons and system stability, but also introduced constraints. In some cases, semantically similar clusters emerged, particularly in weather-related topics, where recurring terms such as “pressure,” “humidity,” “wind” were dispersed across multiple clusters. This effect may be amplified in datasets containing unfiltered automated “weather bot” content. Empty or low-content clusters occasionally appeared due to sparse coverage, noise, or embedding collapse. The spatial distribution of events significantly shaped clustering performance. MultiGraph produced the lowest inter-cluster similarity for geographically concentrated events (Chile wildfires, Napa earthquake), benefiting from modality-specific graph encoding and cross-attention fusion. MonoGraph showed better separation in smaller or noisier AOIs (Ahr Valley floods, Hurricane Harvey), where its unified representation was more robust to diffuse or overlapping discourse. Spatially, MultiGraph generated compact, well-defined hotspots with fewer outliers, aligning clusters closely with localized disaster impacts, an asset for targeted response planning. MonoGraph captured broader spatial patterns, useful for representing large-scale narratives but with a tendency to overestimate geographic extent.
Several limitations merit consideration. First, the framework’s performance is sensitive to the quality and balance of input modalities. Second, the current implementation does not explicitly model temporal dynamics, which constrains the ability to capture and analyze the evolution of topics over time. Third, the use of a fixed number of clusters ($K$) may not reflect the true topical complexity of diverse events, suggesting the need for adaptive or non-parametric clustering strategies. Finally, MultiGraph’s scalability to real-time, high-volume streams is untested. These constraints may hinder applicability to scenarios involving extremely high-frequency data or expansive geographic coverage unless complemented by distributed computing or streaming graph processing frameworks. Despite these constraints, both architectures effectively detect geographically specific and semantically coherent topics that align with real-world hazard footprints, offering practical value for enhancing situational awareness and supporting decision-making in disaster management.
@@ 7 Conclusion
This work presented an unsupervised multimodal graph-based methodology for the semantic and spatial analysis of geo-social media data. By jointly modeling semantic embeddings and geographic coordinates through graph-based architectures, the approach addresses limitations of prior models that treat these modalities in isolation. Two architectural variants, MonoGraph and MultiGraph, were introduced to accommodate differing data characteristics and application needs. A composite loss function integrating contrastive, coherence, and alignment terms was key to learning robust, cluster-aware representations that preserve both local and global structure. Extensive evaluation on four real-world disaster datasets demonstrated consistent improvements over state-of-the-art baselines in topic coherence, spatial compactness, and overall topic quality. The MultiGraph variant, employing modality-specific encoding with attention-based fusion, achieved the strongest performance in most settings, while qualitative analysis of Hurricane Harvey (case study) confirmed the interpretability and geographic relevance of the discovered topics. The proposed methodology offers a scalable, principled approach to integrate symbolic and spatial signals for geo-social media analysis. Although the current implementation focuses on text and location, it can be extended to additional modalities such as sentiment, temporal dynamics, or imagery, enabling richer, more context-aware analyses in domains including disaster management, urban planning, and public health surveillance. Future directions include incorporating time-aware modeling, improving scalability for real-time data streams, and leveraging social network structures to enhance community-level insights. By unifying multimodal signals in an end-to-end learning model, this work advances toward more holistic and interpretable models for understanding the complex information landscapes of social media.

### v2

@@ 7.1 Topic Interpretation
Other topics represent broader or more indirect engagement with the event. Topic 8 in MonoGraph and Topic 5 and in MultiGraph, captured meteorological observations, with keywords such as “temperature,” “humidity,” “clouds,” “rain,” and “wind,” are more evenly distributed across central Texas, including Austin and San Antonio. These spatial patterns suggest contributions from users outside the most severely affected areas, often in the form of commentary or weather updates rather than first-hand impact reports. Weather-related discourse is further reflected in Topics 2, 7, 5 and 9, in MonoGraph and Topics 3, 4, 9, 6 and 8 in MultiGraph, featuring overlapping terms such as “pressure,” “gust,” “humidity,” “wind,” and “clouds.” While this lexical overlap reduces topic distinctiveness, it underscores the recurring presence of weather-related signals across the corpus, a common characteristic in disaster-related social media data. This also illustrates the trade-off of fixing the number of topics ($k_{\text{means}}=15$), which facilitates stability but can lead to partial redundancy. Finally, content-sharing and localized observational behavior emerge as a consistent pattern. For example, certain clusters contain terms such as “posted,” “photo,” and place names like “meyerland,” “katy,” and “downtown,” suggesting the dissemination of real-time visual documentation and location-tagged updates by affected residents. This pattern reflects well-known crisis communication behavior on social media, where users share on-the-ground images to inform others and raise awareness.
@@ 8 Discussion and Limitations
Across all datasets, both MonoGraph and MultiGraph consistently outperform baseline methods, achieving higher intra-cluster cohesion and lower inter-cluster similarity. Their relative strengths vary with event characteristics, reflecting sensitivity to spatial scale, discourse structure, and data quality. Fixing the number of clusters at $k_{\text{means}}=15$ facilitates direct cross-model comparison and stabilizes evaluation, but also introduces constraints. In particular, a fixed $k_{\text{means}}$ may not reflect the true topical complexity of heterogeneous events. In some cases, semantically similar clusters emerge, especially for weather-related topics where recurring terms such as ”pressure,” ”humidity” and ”wind” are distributed across multiple clusters. This effect may be further amplified in datasets containing unfiltered automated or bot-generated content. Occasional empty or low-content clusters also appear in sparsely populated or noisy regions, likely due to limited message coverage or embedding collapse.
From a spatial perspective, event structure strongly shapes clustering behavior. MultiGraph produces compact, well-defined spatial clusters with fewer outliers in geographically concentrated events (e.g., Chile wildfires and the Napa earthquake), benefiting from modality-specific graph construction and cross-attention fusion. MonoGraph exhibits more robust separation in smaller or noisier areas of interest (e.g., Ahr Valley floods and Hurricane Harvey), where a unified representation better accommodates diffuse or overlapping discourse. As a result, MultiGraph tends to align clusters more closely with localized hazard impacts, while MonoGraph captures broader spatial narratives but may overestimate geographic extent.
Several limitations of the proposed framework naturally motivate directions for future research. Model performance depends on the quality and balance of the input modalities, which may vary across platforms and events and affect generalization. Temporal dynamics are not explicitly modeled, limiting the ability to capture topic evolution and cascading effects over time. While fixing $k_{\text{means}}$ enables controlled comparison across methods, adaptive or stability-driven strategies for determining the number of clusters, such as consensus-based or non-parametric approaches, represent an important extension. In addition, automated or bot-generated content is not explicitly filtered, which may introduce redundancy or noise in certain datasets. Evaluation currently relies on intrinsic topic quality metrics and spatial statistics; incorporating human-centered assessment, such as expert judgment of topic interpretability and relevance, or downstream task-based validation, would further strengthen claims of practical utility. Finally, the scalability of the MultiGraph architecture to real-time, high-volume data streams remains untested and may require distributed or streaming graph-processing solutions, alongside exploration of alternative graph construction strategies, additional modalities (e.g., sentiment or temporal signals), and explicit social network structures to enhance community-level insights.
@@ 9 Conclusion
This work presented an unsupervised multimodal graph-based methodology for the semantic and spatial analysis of geo-social media data. By jointly modeling semantic embeddings and geographic coordinates through graph-based architectures, the approach addresses limitations of prior models that treat these modalities in isolation. Two architectural variants, MonoGraph and MultiGraph, were introduced to accommodate differing data characteristics and application needs. A composite loss function integrating contrastive, coherence, and alignment terms was key to learning robust, cluster-aware representations that preserve both local and global structure. Extensive evaluation on four real-world disaster datasets demonstrated consistent improvements over state-of-the-art baselines in topic coherence, spatial compactness, and overall topic quality. The MultiGraph variant, employing modality-specific encoding with attention-based fusion, achieved the strongest performance in most settings, while qualitative analysis of Hurricane Harvey (case study) confirmed the interpretability and geographic relevance of the discovered topics. The proposed methodology offers a scalable, principled approach to integrate symbolic and spatial signals for geo-social media analysis. Although the current implementation focuses on text and location, it can be extended to additional modalities such as sentiment, temporal dynamics, or imagery, enabling richer, more context-aware analyses in domains including disaster management, urban planning, and public health surveillance. By unifying multimodal semantic and spatial signals within an end-to-end learning framework, this work takes a step toward more holistic and interpretable models for understanding the complex information landscapes of geo-social media.
[…]
Acknowledgements
This research was funded by the Austrian Research Promotion Agency (FFG) through the project MOSAIK (grant number 926200) and by the European Union under the Horizon Europe Research and Innovation Actions programme (grant agreement No. 101093003, HORIZON-CL4-2022-DATA-01-01). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them.